OpenStudy (anonymous):

How do I find the price of a cd if 3 cd + 2 DVD=$67. And 2cd+4dvd=$90 ?

OpenStudy (anonymous):

looks to me like you will have to set up a system of equations and use either the elimination methode or the subtution methode to solve this problem

OpenStudy (anonymous):

zbay is exactly correct. But if you have a graphing calculator, you could set up a matrix.

OpenStudy (anonymous):

I set up a graph but got stuck

OpenStudy (anonymous):

do it this way 3c+2d=67 2c+4d=90

OpenStudy (anonymous):

there is your system of equations

OpenStudy (anonymous):

I tried that too but got stuck

OpenStudy (anonymous):

Ok where did you get stuck at? and what methode were you trying?

OpenStudy (anonymous):

I tried both. I set up a chart but sine the two are not right next to each other on the chart I got stuck.

OpenStudy (anonymous):

Oh it's been a while for me but lets try this together

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

Ok, first we want to pick a varible and make it an oppisite inverse of the other equation i am going to do this by multiplying the top function by -2

OpenStudy (anonymous):

(-2)(3c+2d=67) -6c-4d=-134

OpenStudy (anonymous):

Do u know how to do it using a graph? Btw thanks for all your help.

OpenStudy (anonymous):

-6c-4d=-134 2c+4d=90 No i don't think i know how to do the graph

OpenStudy (anonymous):

Ok that is fine

OpenStudy (anonymous):

when we combine them we get (-4c=-44)

OpenStudy (anonymous):

d= 11

OpenStudy (anonymous):

Your awesome ,,, thanks

OpenStudy (anonymous):

c=15

OpenStudy (anonymous):

ok after you have your d value plug it into your orginal equation and solve for c

OpenStudy (anonymous):

Got it:)